How Do I Graph These Equations?
3x+y=5
-2x+3y=4

Found 2 solutions by wuwei96815, sgt_spike:
Answer by wuwei96815(245)

The easiest way to graph an equation is to find the x and y intercepts. You only need two points to graph a linear equation.

3x+y=5
Let x=0
3(0) + y = 5
y=5
You now have the point (0,5) which is called the y-intercept.

3x+y=5
Let y=0
3x + (0) = 5
x=5/3
You now have the point (5/3,0) which is called the x-intercept.

Plot the two points and draw a straight line. Any questions?

















Answer by sgt_spike(2)

First solve for y.
To do this multiply 3x+y=5 by -3 on both sides of the equation.
(-3)3x+y=5(-3)
-9x-3y=-15
subtract this equation from the other equation
-9x-3y=-15
-2x+3y=4
-11x=-11
Divide by -11 on both sides to get x=1
substitute the 1 in the equation -2x+3y=4
-2(1)+3y=4
-2+3y=4
Now solve for y
2-2+3y=4+2
3y=6
y=2
start point (1,2)
Thats how to solve it. DOn't know how to graph it. Sorry